# Perturbation-Robust Predictive Modeling of Social Effects by Network Subspace Generalized Linear Models

### v4

@@ 5.2 Predictive Model Validation
Figure 4 shows the AUC values calculated based on predictions in each individual school by our model and the two benchmarks. The five selected schools in the previous analysis are colored red. It can be seen that overall, the RNC predictions are more accurate, and our model is also more accurate than the standard logistic regression. Due to the significant network effects of the five selected schools, our model provides more accurate predictions. The result shows that overall, the social network effects are sufficiently influential to exhibit differential prediction accuracy.
Since our prediction model interpretations in Table 7 are based on selected schools, we also want to evaluate the effects of this selection procedure. Therefore, we apply the aforementioned 200-fold cross-validation but include the school selection procedure: In each iteration, we first select five schools based on the 199 folds and then focus on model fitting and predicting the hold-out fold constrained within the selected five schools. Note that in this procedure, different schools may be selected for each iteration. Thus, this evaluation also includes the randomness of the selection. The ROC curves aggregated over the 200 folds are shown in Figure 5. The conclusion from Figure 5 is consistent with Figure 4. Both our model and the RNC effectively exploit the network information to achieve more accurate predictions. While the RNC is overall more accurate for prediction, our model’s performance is not far behind.
@@
In summary, our validation experiments show that, whether we consider the selected subset of schools or all of them, the social network effects are strong, and ignoring them results in inferior prediction performance. The RNC provides the best predictive power among the three models, but our model remains competitive. The major drawback of the RNC is the lack of an inference framework, which hinders scientific interpretations of the problem. In contrast, the proposed SP framework offers interpretations, inference, and effective predictions all in one model, with provable robustness to network perturbations.
@@ 6 Discussion
Several interesting directions for expanding our study remain. One particularly intriguing problem is incorporating more general graph neural networks (Scarselli et al., 2008; Kipf and Welling, 2016) into similar subspace models for network-linked data and extending the inference framework to such situations. In a related direction, conformal predictions have been studied for network regression problems (Lunde et al., 2023), but addressing the inference problem with these complications can be more widely useful yet more challenging. Finally, a fundamental problem is using such subspace models to handle spill-over effects of randomized experiments on social networks or even more general causal analysis with network effects (Sinclair et al., 2012; Phan and Airoldi, 2015; Lee and Ogburn, 2021; Hayes et al., 2022), or even other causal analysis involving network-mediated effects. Formulating a spill-over or mediation causal model in the subspace format would be crucial when we want to generalize the proposed framework for such analysis.
@@ A.1 Finalized model on all 25 schools
The estimated parameters and p-values (after variable selection by backward elimination) using the averaged network of two waves in all schools are summarized in Table 8. Under our model, $r$ is detected to be 00, and the $\chi^{2}$ test for the existence of the network effect yields a p-value $<10^{-8}$, suggesting the statistical significance of the network information.
Table 8: Estimated coefficients and p-values of our model, standard logistic regression and RNC using the average network of two waves involving all schools.
@@
Gender:Male -0.443 $<10^{-3}$0.419 $<10^{-3}$0.497
Grade -0.323 $<10^{-3}$0.256 $<10^{-3}$0.359
@@
Home Language 0.272 $<10^{-3}$ 0.324 $<10^{-3}$ 0.241
Treatment 0.837 $<10^{-3}$ 0.820 $<10^{-3}$ 0.800
@@ A.2 Robustness validation with three network constructions
Table 9: Estimated coefficients and p-values of our model using three different versions of the friendship networks. The blanks indicate that the variables are removed in the backward elimination procedure.
@@
Gender:Male -0.443 $<10^{-3}$0.429 $<10^{-3}$0.427 $<10^{-3}$
Grade -0.323 $<10^{-3}$0.252 $<10^{-3}$0.257 $<10^{-3}$
@@
Home Language 0.272 $<10^{-3}$ 0.316 $<10^{-3}$ 0.366 $<10^{-3}$
Treatment 0.837 $<10^{-3}$ 0.842 $<10^{-3}$ 0.851 $<10^{-3}$
@@
To evaluate the robustness of our inference results to network perturbations, we fit our model separately using the Wave I and Wave II networks. In Table 9, we present the results of the same analysis with backward elimination, showing the estimated parameters and p-values for the models based on the three versions of the networks. Despite the significant variations in the networks at the edge level (Figure 1), our method consistently selects the same set of variables, with only minor numerical differences across all three versions. These findings demonstrate the robustness of our framework. As noted in Le and Li [2022], this stability arises because, although the individual edges in the friendship networks experienced substantial changes, the overall spectral structure of the adjacency matrices remained stable.
@@ A.3 Predictive comparisons with embedding-based methods
Figure 6: Prediction performance comparison for each school between using the observed network (Wave I + Wave II), and the embedding similarity of DeepWalk, Node2Vec and Diff2Vec.
@@ A.4 Covariate correlation with network structures in the refined analysis
Figure 7 displays the gender, grade and race information on the other four selected schools in our refined analysis. These covariates displays evident cohesive pattern based on the network structure, which explains why the inference can be different for them between our model and the logistic regression without using the network information.
@@ Appendix B Proofs for theoretical results
$\max\left\{\big{\|}\hat{\mathcal{P}}_{R}-\mathcal{P}_{R}\big{\|},\big{\|}\hat{\mathcal{P}}_{C}-\mathcal{P}_{C}\big{\|},\big{\|}\hat{\mathcal{P}}_{N}-\mathcal{P}_{N}\big{\|}\right\}\leq C_{1}\bar{\tau}_{n},$
@@
$\bar{\tau}_{n}=\frac{n^{-3/2}\big{\|}(\tilde{W}\tilde{W}^{T}-WW^{T})Z\big{\|}}{\min\left\{\left(1-\sigma_{r+1}\right)^{3},\sigma_{r+s}^{3}\right\}},$ (25)
@@
These vectors depend on the choices of bases for $\operatorname{col}(X)$, $S_{K}(P)$, and $S_{K}(\hat{P})$ through $Z$, $W$, $\tilde{Z}$ and $\tilde{W}$. Due to the nature of these choices, $g_{i}$ and $\tilde{g}_{i}$ can be approximately aligned through an almost rotation matrix defined by
@@
$\left\|{g}_{i}-T_{n}^{\top}\tilde{g}_{i}\right\|$$\leq$ $C_{2}n^{1/2}\bar{\tau}_{n},$ (27)
$n^{-1}\sum_{i=1}^{n}\left\|{g}_{i}-T_{n}^{\top}\tilde{g}_{i}\right\|$$\leq$ $C_{2}\bar{\tau}_{n},$ (28)
$\left\|T_{n}^{\top}T_{n}-I_{K+p-r}\right\|_{\infty}$$\leq$ $C_{2}\bar{\tau}_{n}.$ (29)
@@
We will prove that each term on the right-hand side of the above inequality is of order $O(n^{1/2}\bar{\tau}_{n})$. For the first term, since columns of $Z_{1:r}$ are of norm $\sqrt{n}$, by Proposition 1,
@@
$\leq\sum_{i=1}^{r}\left\|(\hat{\mathcal{P}}_{R}-\mathcal{P}_{R})Z_{i}\right\|\leq C_{1}rn^{1/2}\bar{\tau}_{n}.$
@@
$\left\|{Z}^{\top}_{i,(r+1):p}-n^{-1}{Z}^{\top}_{(r+1):p}\tilde{Z}_{(r+1):p}\tilde{Z}^{\top}_{i,(r+1):p}\right\|$ $\leq\sum_{i=1}^{p-r}\left\|(\hat{\mathcal{P}}_{C}-\mathcal{P}_{C})Z_{r+i}\right\|\leq C_{1}(p-r)n^{1/2}\bar{\tau}_{n}.$
@@
As before, we will show that each sum on the right-hand side of (B) is of order $O(\bar{\tau}_{n})$. Regarding the first sum, by the Cauchy-Schwartz inequality and Proposition 1,
@@
$\leq n^{1/2}\sum_{i=1}^{r}\left\|(\hat{\mathcal{P}}_{R}-\mathcal{P}_{R})Z_{i}\right\|\leq C_{1}rn\bar{\tau}_{n}.$
Similarly, the second sum and the third sum on the right-hand side of (B) are bounded by $C_{1}(p-r)n\bar{\tau}_{n}$ and $C_{1}(K-r)n\bar{\tau}_{n}$, respectively. These inequalities and (B) then imply (28).
@@
We will show that each diagonal block is of order $O(\bar{\tau}_{n})$. Regarding the first block, for any unit vectors $u,v\in\mathbb{R}^{k-r}$, by Proposition 1 we have
@@
$\leq$ $C_{1}n\bar{\tau}_{n}\|Z_{1:r}u\|\|Z_{1:r}v\|.$
Since columns of $Z_{1:r}$ are of norm $n^{1/2}$, it follows that $\|Z_{1:r}u\|\|Z_{1:r}v\|\leq rn$. Because $u$ and $v$ are arbitrary, this implies the infinity norm of the first diagonal block is at most $C_{1}rn^{2}\bar{\tau}_{n}$. The same argument can be applied to the second and third diagonal blocks to show that their infinity norms are bounded by $C_{1}(p-r)n^{2}\bar{\tau}_{n}$ and $C_{1}(K-r)n^{2}\bar{\tau}_{n}$, respectively. Together, these bounds imply (29) and the proof of Lemma 1 is complete. ∎
@@
Based on Lemma 1, $T_{n}$ is nonsingular when $\bar{\tau}_{n}<1/C_{2}$. Then we try to bound $T_{n}$’s eigenvalue $\lambda$ by bound $\|T_{n}u_{\lambda}\|$, where $\lambda u_{\lambda}=T_{n}u_{\lambda}$ and $u_{\lambda}\in\mathbb{R}^{K+p-r}$ is a unit vector. By infinity norm bound (29) in Lemma 1

### v6

@@ 5.2 Predictive Model Validation
Figure 4 shows the AUC values calculated based on predictions in each individual school by our model and the two benchmarks. The five selected schools in the previous analysis are colored red. The results show that our model consistently outperforms standard logistic regression and RNC, especially in the five schools with strong network effects. This indicates that our model effectively exploits the network information to provide more accurate predictions. The result also shows that overall, the social network effects are sufficiently influential to exhibit differential prediction accuracy.
Since our prediction model interpretations in Table 12 are based on selected schools, we also want to evaluate the effects of this selection procedure. Therefore, we apply the aforementioned 200-fold cross-validation but include the school selection procedure: In each iteration, we first select five schools based on the 199 folds and then focus on model fitting and predicting the hold-out fold constrained within the selected five schools. Note that different schools may be selected for each iteration in this procedure. Thus, this evaluation also includes the randomness of the selection. The ROC curves aggregated over the 200 folds are shown in Figure 5. The conclusion from Figure 5 is consistent with Figure 4.
@@
In summary, our validation experiments show that, whether we consider the selected subset of schools or all of them, the social network effects are strong, and ignoring them results in inferior prediction performance. Our model provides the best predictive power among the three models. Compared to RNC, the proposed framework offers interpretability, valid inference, and accurate predictions, with provable robustness to network perturbations.
@@ 6 Discussion
Several interesting directions for expanding our study remain. One particularly intriguing problem is incorporating more general graph neural networks (Scarselli et al., 2008; Kipf and Welling, 2016) into similar subspace models for network-linked data and extending the inference framework to such situations. In a related direction, conformal predictions have been studied for network regression problems (Lunde et al., 2023), but adapting a formal inference framework to handle these additional complications would be both more widely useful and more challenging. Finally, a fundamental problem is using such subspace models to handle spill-over effects of randomized experiments on social networks or even more general causal analysis with network effects (Sinclair et al., 2012; Phan and Airoldi, 2015; Lee and Ogburn, 2021; Hayes et al., 2022), or even other causal analysis involving network-mediated effects. Formulating a spill-over or mediation causal model in the subspace format would be a crucial step for generalizing the proposed framework for such analyses.
@@ A.1 Finalized model on all 25 schools
The estimated parameters and $p$-values before and after variable selection by backward elimination using the averaged network of two waves in all schools are summarized in Tables 15 and 16. Under our model, $r$ is detected to be 0, and the $\chi^{2}$ test for the existence of the network effect yields a $p$-value $10^{-8}$, suggesting the statistical significance of the network information.
Table 15: Estimated coefficients and $p$-values (before variable selection through backward elimination) of our model, standard logistic regression and RNC using the average network of two waves involving all schools.
@@
Gender: Male -0.417 $10^{-3} -$0.405 $10^{-3} -$0.497
Grade -0.270 $10^{-3} -$0.271 $10^{-3} -$0.359
Friends like house 0.100 0.059 0.102 0.106 0.051
Home language 0.300 $10^{-3}$ 0.305 $10^{-3}$ 0.241
Treatment 0.846 $10^{-3}$ 0.828 $10^{-3}$ 0.800
GPA 0.085 0.066 0.067 0.228 -0.228
Influencer 0.187 0.054 0.234 0.036 0.311
Race: White -0.079 0.218 -0.082 0.416 -0.342
Race: Black 0.087 0.241 0.045 0.715 -0.211
Race: Hispanic -0.093 0.174 -0.113 0.254 -0.209
Race: Asian 0.086 0.293 0.074 0.637 -0.048
School 3 0.668 0.040 0.916 $10^{-3}$
School 10 -0.345 0.195 0.157 0.481
School 13 -0.558 0.068 -0.310 0.214
School 19 -1.326 0.002 -1.140 $10^{-3}$
School 20 -1.874 $10^{-3} -$1.566 $10^{-3}$
School 21 -1.343 0.005 -1.209 $10^{-3}$
School 22 -1.479 0.001 -1.893 $10^{-3}$
School 24 -1.130 0.002 -0.888 $10^{-3}$
School 26 -1.020 0.013 -0.392 0.137
School 29 -1.467 0.002 -0.721 0.006
School 31 -0.278 0.259 -0.048 0.819
School 34 -0.611 0.052 -0.345 0.105
School 35 -0.162 0.340 0.212 0.315
School 40 0.805 0.017 1.047 $10^{-3}$
School 42 -0.412 0.151 -0.243 0.273
School 44 -1.376 0.002 -0.138 0.564
School 45 -1.770 $10^{-3} -$1.497 $10^{-3}$
School 48 -0.098 0.406 0.112 0.617
School 49 -0.354 0.181 0.193 0.365
School 51 -0.294 0.233 0.273 0.178
School 56 -0.237 0.272 -0.187 0.375
School 58 -1.163 0.002 -0.770 $10^{-3}$
School 60 -1.175 0.001 -0.716 0.001
Table 16: Estimated coefficients and $p$-values (after variable selection through backward elimination) of our model, standard logistic regression and RNC using the average network of two waves involving all schools.
Our Model Logistic Reg RNC
coef. $p$-value coef. $p$-value coef. $p$-value
Gender: Male -0.443 $10^{-3} -$0.419 $10^{-3} -$0.497
Grade -0.323 $10^{-3} -$0.256 $10^{-3} -$0.359
@@
Home language 0.272 $10^{-3}$ 0.324 $10^{-3}$ 0.241
Treatment 0.837 $10^{-3}$ 0.820 $10^{-3}$ 0.800
@@
School 3 0.969 $10^{-3}$
School 10 -0.803 $10^{-3}$
School 13 -0.977 $10^{-3}$
School 19 -1.709 $0^{-3} -1.1$22 $10^{-3}$
School 20 -2.363 $10^{-3} -$1.487 $10^{-3}$
School 21 -1.815 $10^{-3} -$1.248 0.001
School 22 -1.885 $10^{-3} -$1.850 $10^{-3}$
School 24 -1.557 $10^{-3} -$0.852 $10^{-3}$
School 26 -1.375 $10^{-3}$
School 29 -1.869 $10^{-3} -$0.702 0.001
School 34 -1.145 $10^{-3} -$0.306 0.024
School 40 1.009 $10^{-3}$
School 42 -0.845 $10^{-3}$
School 44 -1.864 $10^{-3}$
School 45 -2.080 $10^{-3} -$1.457 $10^{-3}$
School 49 -0.777 $10^{-3}$
School 51 -0.724 $10^{-3}$ 0.275 0.022
School 56 -0.619 0.002
School 58 -1.580 $10^{-3} -$0.732 $10^{-3}$
School 60 -1.511 $10^{-3} -$0.612 $10^{-3}$
The RNC columns for school fixed effects are blank because fixed effects are not identifiable under RNC’s penalty. Additionally, we observe that many school effect terms differ significantly between the standard logistic regression and our model. This suggests that the network effect captured by our model better explains the heterogeneity within schools.
@@ A.2 Robustness validation with three network constructions
Table 17: Estimated coefficients and $p$-values of our model using three versions of the friendship network (Wave I, Wave II, and Wave I-II average). The blanks indicate that the variables are removed in the backward elimination procedure.
@@
Gender: Male -0.443 $10^{-3} -$0.429 $10^{-3} -$0.427 $10^{-3}$
Grade -0.323 $10^{-3} -$0.252 $10^{-3} -$0.257 $10^{-3}$
@@
Home language 0.272 $10^{-3}$ 0.316 $10^{-3}$ 0.366 $10^{-3}$
Treatment 0.837 $10^{-3}$ 0.842 $10^{-3}$ 0.851 $10^{-3}$
@@
School 3 1.258 $10^{-3}$
School 10 -0.803 $10^{-3} -$0.858 $10^{-3}$
School 13 -0.977 $10^{-3} -$1.076 $10^{-3}$
School 19 -1.709 $10^{-3} -$1.779 $10^{-3}$
School 20 -2.363 $10^{-3} -$2.324 $10^{-3} -$1.283 $10^{-3}$
School 21 -1.815 $10^{-3} -$2.033 $10^{-3}$
School 22 -1.885 $10^{-3} -$2.331 $10^{-3}$
School 24 -1.557 $10^{-3} -$1.682 $10^{-3} -$0.616 0.002
School 26 -1.375 $10^{-3} -$1.429 $10^{-3}$
School 27 -0.766 $10^{-3}$ 0.882 $10^{-3}$
School 29 -1.869 $10^{-3} -$1.543 $10^{-3}$
School 31 -0.748 $10^{-3}$
School 34 -1.145 $0^{-3} -1.1$57 $10^{-3}$
School 35 -0.909 $10^{-3}$
School 40 1.281 $10^{-3}$
School 42 -0.845 $10^{-3} -$0.936 $10^{-3}$
School 44 -1.864 $10^{-3} -$0.968 $10^{-3}$
School 45 -2.080 $10^{-3} -$2.304 $0^{-3} -1.1$65 $10^{-3}$
School 48 -0.726 $10^{-3}$
School 49 -0.777 $10^{-3} -$0.843 $10^{-3}$
School 51 -0.724 $10^{-3}$ 0.729 $10^{-3}$
School 56 -0.619 0.002 -0.936 $10^{-3}$
School 58 -1.580 $10^{-3} -$1.627 $10^{-3}$
School 60 -1.511 $10^{-3} -$1.622 $10^{-3}$
To evaluate the robustness of our inference results to network perturbations, we fit our model separately using the Wave I and Wave II networks. In Table 17, we present the results of the same analysis with backward elimination, showing the estimated parameters and $p$-values for the models based on the three versions of the networks. Despite substantial edge‑level variation between waves (Fig. 1), the selected variables are identical (up to minor numeric differences). These findings demonstrate the robustness of our framework. As noted in Le and Li [2022], this stability arises because, although the individual edges in the friendship networks experienced substantial changes, the overall spectral structure of the adjacency matrices remained stable.
@@ A.3 Predictive comparisons with embedding-based methods
Figure 6: Prediction performance comparison for each school between the observed network (Wave I + Wave II), and the embedding similarity of DeepWalk, Node2Vec and Diff2Vec.
@@ A.4 Covariate correlation with network structures in the refined analysis
Figure 7 displays the gender, grade and race information for the other four selected schools in our refined analysis. These covariates display a cohesive pattern based on the network structure, which explains why inference can differ between our model and logistic regression without network information.
@@ Appendix B Proofs for theoretical results
$\max\left\{\big{\|}\hat{\mathcal{P}}_{R}-\mathcal{P}_{R}\big{\|},\big{\|}\hat{\mathcal{P}}_{C}-\mathcal{P}_{C}\big{\|},\big{\|}\hat{\mathcal{P}}_{N}-\mathcal{P}_{N}\big{\|}\right\}\leq C_{1}\tau_{n},$
@@
$\tau_{n}=\frac{n^{-3/2}\big{\|}(\tilde{W}\tilde{W}^{\top}-WW^{\top})Z\big{\|}}{\min\left\{\left(1-\sigma_{r+1}\right)^{3},\sigma_{r+s}^{3}\right\}},$ (25)
@@
[…]
@@
$\left\|{g}_{i}-T_{n}^{\top}\tilde{g}_{i}\right\|$ $\leq$ $C_{2}n^{1/2}\tau_{n},$ (27)
$n^{-1}\sum_{i=1}^{n}\left\|{g}_{i}-T_{n}^{\top}\tilde{g}_{i}\right\|$ $\leq$ $C_{2}\tau_{n},$ (28)
$\left\|T_{n}^{\top}T_{n}-I_{K+p-r}\right\|_{\infty}$ $\leq$ $C_{2}\tau_{n}.$ (29)
@@
We will prove that each term on the right-hand side of the above inequality is of order $O(n^{1/2}\tau_{n})$. For the first term, since columns of $Z_{1:r}$ are of norm $\sqrt{n}$, by Proposition 1,
@@
$\leq\sum_{i=1}^{r}\left\|(\hat{\mathcal{P}}_{R}-\mathcal{P}_{R})Z_{i}\right\|\leq C_{1}rn^{1/2}\tau_{n}.$
@@
$\left\|{Z}^{\top}_{i,(r+1):p}-n^{-1}{Z}^{\top}_{(r+1):p}\tilde{Z}_{(r+1):p}\tilde{Z}^{\top}_{i,(r+1):p}\right\|$ $\leq\sum_{i=1}^{p-r}\left\|(\hat{\mathcal{P}}_{C}-\mathcal{P}_{C})Z_{r+i}\right\|\leq C_{1}(p-r)n^{1/2}\tau_{n}.$
@@
As before, we will show that each sum on the right-hand side of (B) is of order $O(\tau_{n})$. Regarding the first sum, by the Cauchy-Schwartz inequality and Proposition 1,
@@
$\leq n^{1/2}\sum_{i=1}^{r}\left\|(\hat{\mathcal{P}}_{R}-\mathcal{P}_{R})Z_{i}\right\|\leq C_{1}rn\tau_{n}.$
Similarly, the second sum and the third sum on the right-hand side of (B) are bounded by $C_{1}(p-r)n\tau_{n}$ and $C_{1}(K-r)n\tau_{n}$, respectively. These inequalities and (B) then imply (28).
@@
We will show that each diagonal block is of order $O(\tau_{n})$. Regarding the first block, for any unit vectors $u,v\in\mathbb{R}^{k-r}$, by Proposition 1 we have
@@
$\leq$ $C_{1}n\tau_{n}\|Z_{1:r}u\|\|Z_{1:r}v\|.$
Since columns of $Z_{1:r}$ are of norm $n^{1/2}$, it follows that $\|Z_{1:r}u\|\|Z_{1:r}v\|\leq rn$. Because $u$ and $v$ are arbitrary, this implies the infinity norm of the first diagonal block is at most $C_{1}rn^{2}\tau_{n}$. The same argument can be applied to the second and third diagonal blocks to show that their infinity norms are bounded by $C_{1}(p-r)n^{2}\tau_{n}$ and $C_{1}(K-r)n^{2}\tau_{n}$, respectively. Together, these bounds imply (29) and the proof of Lemma 1 is complete. ∎
@@
Based on Lemma 1, $T_{n}$ is nonsingular when $\tau_{n}<1/C_{2}$. Then we try to bound $T_{n}$’s eigenvalue $\lambda$ by bound $\|T_{n}u_{\lambda}\|$, where $\lambda u_{\lambda}=T_{n}u_{\lambda}$ and $u_{\lambda}\in\mathbb{R}^{K+p-r}$ is a unit vector. By infinity norm bound (29) in Lemma 1